40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 15 00:40:19 2019, gautam, Update, SUS, ETMY diagnosis plan ETMY_sensorSpectra.pdf
    Reply  Wed May 15 10:57:57 2019, gautam, Update, SUS, EY chamber opened 
       Reply  Wed May 15 17:46:24 2019, gautam, Update, SUS, ETMY inspection firstContactFiber.JPGETMY_sensorSpectra.pdf
          Reply  Wed May 15 19:36:29 2019, Koji, Update, SUS, ETMY instepction 
             Reply  Thu May 16 13:07:14 2019, gautam, Update, SUS, First contact residue removal etmy_sensors.pdfetmy_BRmode.pdf
                Reply  Thu May 16 23:31:55 2019, gautam, Update, SUS, ETMY suspension characterization 6x
                   Reply  Fri May 17 17:01:08 2019, gautam, Update, SUS, ETMY suspension characterization etmy_oplevs.pdf
          Reply  Mon May 20 11:33:46 2019, gautam, Update, SUS, ITMY inspection 
             Reply  Mon May 20 17:12:57 2019, gautam, Update, SUS, ETMY LL adjustment 
                Reply  Mon May 20 22:06:07 2019, gautam, Update, SUS, ITMY also kicked 
                   Reply  Tue May 21 00:15:21 2019, gautam, Update, SUS, Main objectives of vent achieved (?) etmy_oplevs_20190520.pdfETMY_cplxTF.pdfETMY_diagComp.pdfScreen_Shot_2019-05-21_at_12.37.08_AM.png
                      Reply  Tue May 21 21:33:27 2019, gautam, Update, SUS, ETMY HR face cleaned 
                         Reply  Wed May 22 11:53:50 2019, gautam, Update, SUS, ETMY EQ stops backed out ETMY_eigenmodes.pdf
Message ID: 14623     Entry time: Mon May 20 11:33:46 2019     In reply to: 14611     Reply to this: 14625
Author: gautam 
Type: Update 
Category: SUS 
Subject: ITMY inspection 

With Chub providing illumination via the camera viewport, I was able to take photos of ITMY this morning. All the magnets look well clear of the OSEMs, with the possible exception of UR. I will adjust the position of this OSEM slightly. To test if this fix is effective, I will then cycle the bias voltage to the ITM between 0 and the maximum allowed, and check if the optic gets stuck.

ELOG V3.1.3-